# Sharp one-point estimates and Minkowski content for the scaling limit of three-dimensional loop-erased random walk

### v1

@@ 5.2 Off-diagonal two-point estimate
$E\Big{[}J^{m}_{s}(z)J^{m}_{s^{\prime}}(w)\Big{]}\to E\Big{[}J_{s}(z)J_{s^{\prime}}(w)\Big{]}\mbox{ as $m\to\infty$}$ (5.9)
and
$E\Big{[}\big{(}J^{m}_{s}(z)-J_{s^{\prime}}(z)\big{)}\big{(}J^{m}_{s}(w)-J^{m}_{s^{\prime}}(w)\big{)}\Big{]}\leq O_{V}\left(2^{-\delta\sqrt{s}}\right).$ (5.10)
@@
Thanks to Corollary 5.3, we can conduct much of the analysis of $K(z,w,s,s^{\prime})$ in the discrete. Indeed, by Corollary 5.3, to prove Proposition 5.2 it suffices to prove the following estimate for the renormalized indicator function in the discrete space. For the purpose of proving Proposition 5.2, it suffices to work with dyadic scales, since Proposition 2.2 only requires understanding the one and two point Green’s function on exponential scales.
@@
Let $V\in{\cal D}^{o}$ be a dyadic box. Take $s\geq 1$, $s^{\prime}\in[s,s+1]$ and let $z,w\in V$ satisfying $|z-w|\geq 2^{-\sqrt{s}/100}$ (the assumption in Proposition 5.2). Then there exists $M_{2}(s)>0$, such that for all $m=2^{n}\geq M_{2}$ for some $n\in{\mathbb{Z}^{+}}$, one has
@@
$\begin{split}&\Big{|}E\left[J^{m}_{s}(z)\big{(}J^{m}_{s}(w)-J^{m}_{s^{\prime}}(w)\big{)}\right]\Big{|}\\ &\leq\max_{i=0,1}\Big{|}E\left[J^{m}_{s}(z)\big{(}J^{m}_{s_{i}}(w_{m})-J^{m}_{s^{\prime}_{1-i}}(w_{m})\big{)}\right]\Big{|}+O_{V}\left(2^{-cs^{2}}\right).\end{split}$ (5.18)
@@
$A(\widetilde{s})=a_{n}\left(B(2^{-\widetilde{s}})\cap 2^{-n}{\mathbb{Z}}^{3}\right)$
and
$Q=P\big{(}B(z,2^{-s})\cap\eta_{m}\neq\emptyset,w_{m}\in\eta_{m}\big{)}$
@@
$\begin{split}&\Big{|}E\left[J^{m}_{s}(z)\big{(}J^{m}_{s_{i}}(w_{m})-J^{m}_{s^{\prime}_{1-i}}(w_{m})\big{)}\right]\Big{|}\\ \leq\;&2^{(3-\beta)s}\left[\Big{(}1+O\big{(}2^{-\sqrt{s}}\big{)}\Big{)}A(s_{i})2^{(3-\beta)s_{i}}-\Big{(}1+O\big{(}2^{-\sqrt{s}}\big{)}\Big{)}A(s^{\prime}_{1-i})2^{(3-\beta)s^{\prime}_{1-i}}\right]Q\\ \leq\;&\Big{[}{\rm I}+O_{V}\big{(}2^{-\sqrt{s}}\big{)}{\rm II}\Big{]}O_{V}\Big{(}2^{(3-\beta)(\sqrt{s}/100-n)}\Big{)}.\end{split}$ (5.21)
@@
By one-point estimates (2.27) and (2.25),
@@
$g(z,w)$ $\coloneqq\lim_{r\to 0}\mathbb{E}\left(J_{r}(z)J_{r}(w)\right)$
$=\lim_{r\to 0}\left(r^{2(3-\beta)}P\left({\cal K}\cap B(z,r)\neq\emptyset,\,{\cal K}\cap B(w,r)\neq\emptyset\right)\right.$ (5.24)
exists for each $z,w\in\Omega$. This is because, starting from (5.2), it is easy to see that for $m\in\mathbb{Z}$, $J_{m}(V)$ has an $L^{2}$- and a.s.-limit (the latter requiring the use of Chebyshev inequality and Borel-Cantelli lemma), which is also the limit if we let $r$ tend to 00 continuously. See the proof of [15, Theorem 3.1] for a complete argument.
Recall that $\hat{x}=(1/2,0,0)$ is a reference point and that for any $z,w\in\Omega$, $d_{z}=\min\{|z|,1-|z|\}$ and $d_{z,w}=\min\{d_{z},d_{w},|z-w|\}$. Given $m\in\mathbb{R}^{+}$, $\eta_{m}$ denotes the LERW on $m^{-1}\mathbb{Z}^{3}$. Let us fix $z,w\in\Omega$ with $d_{z,w}>e^{-s^{2}}$ and take $0<r<e^{-s^{4}}$. The asymptotic independence from Proposition 2.14 implies that there exists $M_{2}(r,d_{z,w},d_{z},d_{w})>0$ such that for all $m\geq M_{2}$
$\frac{P\left(\begin{array}[]{c}\eta_{m}\cap B(z_{m},r)\neq\emptyset,\\ \,\eta_{m}\cap B(w_{m},r)\neq\emptyset\end{array}\right)}{P\left(z_{m},w_{m}\in\eta_{m}\right)}=\left(\frac{P\left(\eta_{m}\cap B(\hat{x},r)\neq\emptyset\right)}{P(\hat{x}\in\eta_{m})}\right)^{2}\Big{[}1+O\big{(}d_{z,w}\big{)}\Big{]}.$ (5.25)
From (5.25) and Propositions 4.5 and 3.1, we can strengthen the convergence in (5.24) as in the proof of Theorem 1.4, so that
$P\left({\cal K}\cap B(z,r)\neq\emptyset,{\cal K}\cap B(w,r)\neq\emptyset\right)=c_{1}g(z,w)r^{2(3-\beta)}\Big{[}1+O\big{(}d_{z,w}^{-c}r^{\delta}\big{)}\Big{]};$ (5.26)
@@
$P(\eta_{m}\cap B(z,r)\neq\emptyset,\eta_{m}\cap B(w,s)\neq\emptyset)=P(\eta_{m}\cap B(z,r)\neq\emptyset)+P(\eta_{m}\cap B(w,s)\neq\emptyset)\\ -P(\eta_{m}\cap(B(z,r)\cup B(w,r))\neq\emptyset).$ (5.27)
Similarly for $P(K\cap B(z,r)\neq\emptyset,K\cap B(w,r)\neq\emptyset)$. Each one of the terms on the right-hand side of (5.27) approximates the intersection probability of a ball and the scaling limit of the LERW. Proposition 2.6 implies that there exist $c>0$ and $\delta>0$ so that for $r\in(0,e^{-s^{4}})$
@@
$P\left(\eta_{m}\cap B(w,r)\neq\emptyset\right)$ $\simeq P\left({\cal K}\cap B(w,r)\neq\emptyset\right)$
$P\left(\eta_{m}\cap(B(z,r)\cup B(w,r))\neq\emptyset\right)$ $\simeq P\left({\cal K}\cap(B(z,r)\cup B(w,r))\neq\emptyset\right).$
These latter estimates, together with (5.27), imply that for $c>0$ and $\delta>0$ and $r\in(0,e^{-s^{4}})$
$P(\eta_{m}\cap B(z,r)\neq\emptyset,\eta_{m}\cap B(w,s)\neq\emptyset)=\\ P({\cal K}\cap B(z,r)\neq\emptyset,{\cal K}\cap B(w,s)\neq\emptyset)\Big{[}1+O\big{(}d^{-c}_{z,w}m^{-\delta}\big{)}\Big{]}\text{ as }m\to\infty.$ (5.28)
Now, the asymptotic above and (5.26) imply that there exist constants $c^{\prime},\delta^{\prime}>0$ so that
$P(\eta_{m}\cap B(z,r)\neq\emptyset,\eta_{m}\cap B(w,r)\neq\emptyset)=c_{1}g(z,w)r^{2(3-\beta)}\Big{[}1+O\big{(}d_{z,w}^{-c^{\prime}}r^{\delta^{\prime}}\big{)}\Big{]}.$ (5.29)
@@
$\frac{c_{1}g(z,w)}{P\left(z_{m},w_{m}\in\eta_{m}\right)}=\Big{[}1+O\big{(}d_{z,w}^{-c^{\prime\prime}}r^{\delta^{\prime}}\big{)}\Big{]}\left(\frac{P(\eta_{m}\cap B(\hat{x},r)\neq\emptyset)}{r^{3-\beta}P(\hat{x}\in\eta_{m})}\right)^{2}.$ (5.30)
@@
6 Proof of main theorem
In this section, we prove Theorem 1.1.
@@ 6.1 Existence of Minkowski content
To confirm the existence of the Minkowski content, we apply Proposition 2.2 to ${\cal K}$. Consider a fixed $V\in{\cal D}^{o}$ and let
$J_{s}(z)\coloneqq 2^{(3-\beta)s}1_{z\in B({\cal K},2^{-s})}$ (6.1)
be the renormalized indicator function of the blow-up of ${\cal K}$. Write
[…]
@@
By Proposition 2.2, it suffices to verify Conditions 1 and 2.
We note that claims on one-point estimates are easy corollaries of Theorem 1.4. The existence of the two-point Green’s function follows from Theorem 1.6 and its upper bound from (2.31). ∎
@@ 6.2 Relating Minkowski content with limiting occupation measure
$c_{0}(n,s)\coloneqq\frac{E[J^{2^{n}}_{s}(x+\square)]}{2^{n\beta}P(\hat{x}\in\eta_{2^{n}})}.$
@@
The claim (6.2) is an easy corollary of the definition of $J$, Proposition 3.1 and Theorem 1.3, while (6.3) then follows from Theorem 1.4. ∎
@@
[…]
To relate the content with the limiting occupation measure, as in last subsection, again we pass the analysis to the discrete and make use of the asymptotic independence for 3D LERW.
@@
By [20, Proposition 9.1], $V$ is a.s. a continuity set for $\mu$, hence we have
[…]
@@
${\mathbb{P}}(|J_{V}-J_{s}(V)|>\varepsilon/8)\leq C/8.$ (6.8)
@@
$s^{\prime\prime}>s^{\prime}>\max(s_{0},s_{1})\quad\mbox{ and }\quad e^{(3-\beta)(s-s^{\prime})},e^{(3-\beta)(s^{\prime\prime}-s)}<1+\varepsilon_{1}.$
@@
Finally, we see that the inequalities (6.7), (6.14) and (6.9) combined contradict (6.5)!
This finishes the proof of (6.4). ∎
@@
Write $M_{2}\coloneqq M_{2}(V)={\mathbb{E}}[\mu^{2}_{2^{n}}(V)].$ Up-to-constant estimates for two-point functions of 3D LERW (see Proposition 2.13) implies that $M_{2}<\infty$. We fix some $k_{1}=k_{1}(k,\varepsilon,C,V)\in{\mathbb{Z}^{+}}$ such that
@@
For $s>s_{1}$ and $n>n_{3}$, we apply Proposition 2.14 with $V_{i}$ for each $i=1,\ldots,g(k_{1})$ similarly as how we apply [20, Proposition 5.2] to prove [20, Proposition 5.1], and obtain that there exists some universal $C_{2}>0$ such that for all $n>k_{1}^{4})$,
${\mathbb{E}}\Big{[}\big{(}J^{2^{n}}_{s}(V_{i})-c_{0}(n,s)\mu_{2^{n}}(V_{i})\big{)}^{2}\Big{]}=C_{2}2^{-k_{1}^{2}}{\mathbb{E}}[\mu^{2}_{2^{n}}(V_{i})].$ (6.16)
@@
$\begin{split}&\;{\mathbb{P}}\Big{(}\big{|}J_{s}^{2^{n}}(V)-c_{0}\mu_{2^{n}}(V)\big{|}\geq\varepsilon/8\Big{)}\\ \leq&\;{\mathbb{P}}\Big{(}\big{|}J_{s}^{2^{n}}(V)-c_{0}(n,s)\mu_{2^{n}}(V)\big{|}\geq\varepsilon/16\Big{)}+{\mathbb{P}}[\mu_{2^{n}}(V)>M_{1}]\leq C/4.\end{split}$ (6.19)
This finishes the proof of (6.9) ∎

### v2

@@ 5.2 Off-diagonal two-point estimate
$E\Big[J^{m}_{s}(z)J^{m}_{s^{\prime}}(w)\Big]\to E\Big[J_{s}(z)J_{s^{\prime}}(w)\Big]\mbox{ as $m\to\infty$}.$ (5.9)
@@
Thanks to Corollary 5.3, we can conduct much of the analysis of $K(z,w,s,s^{\prime})$ in the discrete. Indeed, by Corollary 5.3, to prove Proposition 5.2 it suffices to prove the following estimate for the renormalized indicator function in the discrete space. For the purpose of proving Proposition 5.2, it suffices to work with dyadic scales, since Proposition 2.2 only requires understanding the one and two point Green’s function for a subsequence of scales tending to infinity.
@@
Let $V\in{\cal D}^{o}$ be a dyadic box. Take $\geq 1, s[s$$^{\prime}\in[s,s+1$] and let $z,w\in V$ satisfying $|z-w|\geq 2^{-\sqrt{s}/100}$ (the assumption in Proposition 5.2). Then there exists $M_{2}(s)>0$, such that for all $m=2^{n}\geq M_{2}$ for some $n\in{\mathbb{Z}^{+}}$, one has
@@
$\Big|E\left[J^{m}_{s}(z)\big(J^{m}_{s}(w)-J^{m}_{s^{\prime}}(w)\big)\right]\Big|$
$\leq\max_{i=0,1}\Big|E\left[J^{m}_{s}(z)\big(J^{m}_{s_{i}}(w_{m})-J^{m}_{s^{\prime}_{1-i}}(w_{m})\big)\right]\Big|+O_{V}\left(2^{-cs^{2}}\right).$ (5.17)
@@
$A(\widetilde{s})=a_{n}\left(B(2^{-\widetilde{s}})\cap 2^{-n}{\mathbb{Z}}^{3}\right)\qquad\mbox{and}\qquad Q=P\big(B(z,2^{-s})\cap\eta_{m}\neq\emptyset,w_{m}\in\eta_{m}\big)$
@@
$\begin{split}&\Big|E\left[J^{m}_{s}(z)\big(J^{m}_{s_{i}}(w_{m})-J^{m}_{s^{\prime}_{1-i}}(w_{m})\big)\right]\Big|\\ \leq\;&2^{(3-\beta)s}\left[\Big(1+O\big(2^{-\sqrt{s}}\big)\Big)A(s_{i})2^{(3-\beta)s_{i}}-\Big(1+O\big(2^{-\sqrt{s}}\big)\Big)A(s^{\prime}_{1-i})2^{(3-\beta)s^{\prime}_{1-i}}\right]Q\\ \leq\;&\Big[{\rm I}+O_{V}\big(2^{-\sqrt{s}}\big){\rm II}\Big]O_{V}\Big(2^{(3-\beta)(\sqrt{s}/100-n)}\Big),\end{split}$ (5.20)
@@
By one-point estimates (2.27) and (2.25), we have
$A(s_{i})\asymp\frac{2^{-(3-\beta)s_{i}}}{2^{-(3-\beta)n}}\ \ \ \text{ and }\ \ \ A(s_{1-i}^{\prime})\asymp\frac{2^{-(3-\beta)s_{1-i}^{\prime}}}{2^{-(3-\beta)n}}.$
Therefore, since $0<\frac{3-\beta}{100}<\frac{3}{100}$, it follows that
@@
$g(z,w)$ $\coloneqq\lim_{s\to\infty}\mathbb{E}\left(J_{s}(z)J_{s}(w)\right)$
$=\lim_{s\to\infty}\left(2^{2(3-\beta)s}P\left({\cal K}\cap B(z,2^{-s})\neq\emptyset,\,{\cal K}\cap B(w,2^{-s})\neq\emptyset\right)\right)$ (5.23)
exists for each $z,w\in\Omega$. This is because, starting from (5.2), it is easy to see that for $m\in\mathbb{Z}$, $J_{m}(V)$ has an $L^{2}$- and a.s.-limit (the latter requiring the use of Chebyshev inequality and Borel-Cantelli lemma), which is also the limit if we let $r$ tend to 0 continuously. See the proof of [16, Theorem 3.1] for a complete argument.
Recall that $\hat{x}=(1/2,0,0)$ is a reference point and that for any $z,w\in\Omega$, $d_{z}=\min\{|z|,1-|z|\}$ and $d_{z,w}=\min\{d_{z},d_{w},|z-w|\}$. Let us fix $z,w\in\Omega$ with $d_{z,w}>e^{-s^{2}}$ and take $0<r<e^{-s^{4}}$. The asymptotic independence from Proposition 2.14 implies that there exists $M_{2}(r,d_{z,w},d_{z},d_{w})>0$ such that for all $m\geq M_{2}$
$\frac{P\big(\eta_{m}\cap B(z_{m},r)\neq\emptyset,\,\eta_{m}\cap B(w_{m},r)\neq\emptyset\big)}{P\big(z_{m},w_{m}\in\eta_{m}\big)}$
$=\left(\frac{P\big(\eta_{m}\cap B(\hat{x},r)\neq\emptyset\big)}{P\big(\hat{x}\in\eta_{m}\big)}\right)^{2}\Big[1+O\big(d_{z,w}^{-c}r^{\delta}\big)\Big].$ (5.24)
Proposition 3.1 ensures that the numerator of the LHS of (5.3) can be accurately approximated by
$P\Big({\cal K}\cap B(z,r)\neq\emptyset,{\cal K}\cap B(w,r)\neq\emptyset\Big).$
Replacing the numerator by the hitting probability for $\cal K$} above, by using Proposition 4.5 and the convergence in (5.3), we see that the quantity
$\frac{P\big(z_{m},w_{m}\in\eta_{m}\big)}{P\big(\hat{x}\in\eta_{m}\big)^{2}}$
converges as $m\to\infty$. Since this quantity does not depend on $r$, we can strengthen the convergence in (5.3) as in the proof of Theorem 1.4, so that
$P\Big({\cal K}\cap B(z,r)\neq\emptyset,{\cal K}\cap B(w,r)\neq\emptyset\Big)=g(z,w)r^{2(3-\beta)}\Big[1+O\big(d_{z,w}^{-c}r^{\delta}\big)\Big],$ (5.25)
@@
$P\Big(\eta_{m}\cap B(z,r)\neq\emptyset,\eta_{m}\cap B(w,s)\neq\emptyset\Big)\\ =P\Big(\eta_{m}\cap B(z,r)\neq\emptyset\Big)+P\Big(\eta_{m}\cap B(w,s)\neq\emptyset\Big)\\ -P\Big(\eta_{m}\cap(B(z,r)\cup B(w,r))\neq\emptyset\Big).$ (5.26)
Similarly for $P({\cal K}\cap B(z,r)\neq\emptyset,{\cal K}\cap B(w,r)\neq\emptyset)$. Each one of the terms on the right-hand side of (5.26) approximates the intersection probability of a ball and the scaling limit of the LERW. Proposition 2.6 implies that there exist $c>0$ and $\delta>0$ so that for $r\in(0,e^{-s^{4}})$
@@
$P\Big(\eta_{m}\cap B(w,r)\neq\emptyset\Big)$ $\simeq P\Big({\cal K}\cap B(w,r)\neq\emptyset\Big);$
$P\Big(\eta_{m}\cap(B(z,r)\cup B(w,r))\neq\emptyset\Big)$ $\simeq P\Big({\cal K}\cap(B(z,r)\cup B(w,r))\neq\emptyset\Big).$
These latter estimates, together with (5.26), imply that for $c>0$ and $\delta>0$ and $r\in(0,e^{-s^{4}})$,
$\begin{split}&P\Big(\eta_{m}\cap B(z,r)\neq\emptyset,\eta_{m}\cap B(w,r)\neq\emptyset\Big)\\ =\;&P\Big({\cal K}\cap B(z,r)\neq\emptyset,{\cal K}\cap B(w,r)\neq\emptyset\Big)\Big[1+O\big(d^{-c}_{z,w}m^{-\delta}\big)\Big]\text{ as }m\to\infty.\end{split}$ (5.27)
Now, (5.27) and (5.25) imply that there exist constants $c^{\prime},\delta^{\prime}>0$ so that
$P\Big(\eta_{m}\cap B(z,r)\neq\emptyset,\eta_{m}\cap B(w,r)\neq\emptyset\Big)=g(z,w)r^{2(3-\beta)}\Big[1+O\big(d_{z,w}^{-c^{\prime}}r^{\delta^{\prime}}\big)\Big].$ (5.28)
@@
$\frac{g(z,w)}{P\Big(z_{m},w_{m}\in\eta_{m}\Big)}=\Big[1+O\big(d_{z,w}^{-c^{\prime\prime}}r^{\delta^{\prime}}\big)\Big]\left(\frac{P\Big(\eta_{m}\cap B(\hat{x},r)\neq\emptyset\Big)}{r^{3-\beta}P\Big(\hat{x}\in\eta_{m}\Big)}\right)^{2}.$ (5.29)
@@
6 Proof of Theorem 1.1.
@@ 6.1 Existence of Minkowski content
To confirm the existence of the Minkowski content, we apply Proposition 2.2 to $\cal K$}. Consider a fixed $V\in{\cal D}^{o}$ and recall $J_{s}(\cdot)$ the renormalized indicator function of the blow-up of $\cal K$}. Write
$J_{s}(V)=\int_{V}J_{s}(z)dz$
@@
By Proposition 2.2, it suffices to verify Conditions 1 and 2 above.
We note that claims on one-point estimates are easy corollaries of Theorem 1.4. The existence of the two-point Green’s function follows from Theorem 1.6 and its upper bound from (2.31). Since we know the rate of convergence for the two-point function (see RHS of (1.15)), the second condition is satisfied. ∎
@@ 6.2 Relating Minkowski content with limiting occupation measure
$c_{0}(n,s)\coloneqq\frac{E[J^{2^{n}}_{s}(\hat{x}+\square)]}{2^{n\beta}P(\hat{x}\in\eta_{2^{n}})}.$
@@
The claim (6.1) is an easy corollary of the definition of $J$, Proposition 3.1 and Theorem 1.3, while (6.2) follows from Theorem 1.4. ∎
@@
$J_{V}=c_{0}\mu(V)\quad\mbox{ a.s.}$ (6.3)
To relate the content to the limiting occupation measure, as in last subsection, again we pass the analysis to the discrete and make use of the asymptotic independence for 3D LERW.
@@
By [22, Proposition 9.4], $V$ is a.s. a continuity set for $\mu$, hence we have
$\mu_{2^{n}}(V)\to\mu(V)\quad{\mathbb{P}}\mbox{-a.s.}$ (6.5)
@@
${\mathbb{P}}\left(|J_{V}-J_{s}(V)|>\varepsilon/8\right)\leq C/8.$ (6.7)
@@
$s^{\prime\prime}>s^{\prime}>\max(s_{0},s_{1})\quad\mbox{ and }\quad\max\big\{e^{(3-\beta)(s-s^{\prime})},e^{(3-\beta)(s^{\prime\prime}-s)}\big\}<1+\varepsilon_{1}.$
@@
Combining (6.6), (6.13) and (6.8), the probability of the event that
$|J_{V}-c_{0}\mu(V)|\leq\frac{5\varepsilon}{8}$
is bigger than $1-\frac{7C}{16}$, which contradicts (6.4)! We hence finish the proof of (6.3). ∎
@@
Write $M_{2}\coloneqq M_{2}(V)=\sup_{n}{\mathbb{E}}[\mu^{2}_{2^{n}}(V)]$. Up-to-constant estimates for two-point functions of 3D LERW (see Proposition 2.13) implies that $M_{2}<\infty$. We fix some $k_{1}=k_{1}(k,\varepsilon,C,V)\in{\mathbb{Z}^{+}}$ such that
@@
For $s>s_{1}$ and $n>n_{3}$, we apply Proposition 2.14 with $V_{i}$ for each $i=1,\ldots,g(k_{1})$ similarly as how we apply [22, Proposition 6.2] to prove [22, Proposition 6.1], and obtain that there exists some universal $C_{2}>0$ such that for all $n>k_{1}^{4})$,
${\mathbb{E}}\Big[\big(J^{2^{n}}_{s}(V_{i})-c_{0}(n,s)\mu_{2^{n}}(V_{i})\big)^{2}\Big]\leq C_{2}2^{-k_{1}^{2}}{\mathbb{E}}[\mu^{2}_{2^{n}}(V_{i})].$ (6.15)
@@
${\mathbb{P}}\Big(\big|J_{s}^{2^{n}}(V)-c_{0}\mu_{2^{n}}(V)\big|\geq\varepsilon/8\Big)$
$\leq{\mathbb{P}}\Big(\big|J_{s}^{2^{n}}(V)-c_{0}(n,s)\mu_{2^{n}}(V)\big|\geq\varepsilon/16\Big)+{\mathbb{P}}[\mu_{2^{n}}(V)>M_{1}]\leq C/4.$
This finishes the proof of (6.8). ∎